Two angles are complementary. If one angle measures 4x+3 and the other angle measures 2x+21, find the measure of each angle.
Answer by stanbon(75887)

Two angles are complementary. If one angle measures 4x+3 and the other angle measures 2x+21, find the measure of each angle.
----
1st angle: 4x+3
2nd angle: 2x+21
------
Equation:
4x+3 + 2x+21 = 90 degrees
6x + 24 = 90
x + 4 = 15
x = 11 degrees
----
1st angle: 4x+3 = 47 degrees
2nd angle: 2x+21 = 43 degrees
===================
Cheers,
Stan H.
====================